Allyson plays for the cook county regoinal football league whuch has ten teams. DURING A SEASON, EACH TEAM PLAYS THE OTHER TEAM TWICE, ONCE AT HOME AND ONE AWAY.HOW MANY TOTAL GAMES ARE PLAYED IN ONE SEASON?: Allyson plays for the cook county regoinal football league whuch has ten teams. DURING A SEASON, EACH TEAM PLAYS THE OTHER TEAM TWICE, ONCE AT HOME AND ONE AWAY.HOW MANY TOTAL GAMES ARE PLAYED IN ONE SEASON?
Answer by stanbon(18994)

It takes two teams to play a game.
Number of pairs of ten is 10C2 = 10*9/1/2 = 45 games
They do that twice, so 90 games would be required for
each team to play the other teams twice..
-------------------------
Cheers,
Stan H.
Question 155027: Allyson plays for the cook county regoinal football league whuch has ten teams. DURING A SEASON, EACH TEAM PLAYS THE OTHER TEAM TWICE, ONCE AT HOME AND ONE AWAY.HOW MANY TOTAL GAMES ARE PLAYED IN ONE SEASON?: Allyson plays for the cook county regoinal football league whuch has ten teams. DURING A SEASON, EACH TEAM PLAYS THE OTHER TEAM TWICE, ONCE AT HOME AND ONE AWAY.HOW MANY TOTAL GAMES ARE PLAYED IN ONE SEASON?
Answer by vleith(1172)

Since there are 10 teams, each week there will be 5 games.
It will take 9 weeks to play every other team once. So it takes 18 weeks to play every team twice (once home and once away).
5*9*2 = 90 games
Using mathematics for permutations is 10!/(10-2)! = 10!/8! = 10*9 = 90
See this video for more info http://collabvsl.wetpaint.com/page/Basic+Probability